A bottle of fruit jucie contains 20% water. How much water must be added to this bottle to produce 8 L of fruit juice that is 50% water
1
Expert's answer
2011-04-07T05:49:03-0400
Let's construct the system of equations.
Denote the initial volume of juice as X, thus it contains 0.2X of the water,& the volume of the water to be added as Y,
(0.2X +Y) = 0.5*8 = 4 ; X +Y =8.
Y = 8 - X,
0.2X + 8 - X = 4
X& = 5;& Y = 3.
Answer: 3 liters of the water must be added.